Lens combinations. You have two lenses - a converg
Answer #1

f1 = 3 cm
f2 = -3 cm

1/f = 1/f1 + 1/f2
1/f = 1/3 - 1/3
f = infinity

Focal length of combination = infinity

Power = 1/f = 0

This combination will behave as simple surface and it will not affect direction of light

I am allowed to answer only 1 question at a time.
I have answered all parts of 1st question